# A Lost Opportunity for Vision-Language Models: A Comparative Study of Online Test-time Adaptation for Vision-Language Models

### v1

@@ Abstract
In the realm of deep learning, maintaining model robustness against distribution shifts is critical. This paper investigates test-time adaptation strategies for vision-language models, with a specific focus on CLIP [37] and its variants. Through a systematic exploration of prompt-based techniques and existing test-time adaptation methods, the study aims to enhance the adaptability and robustness of vision-language models in diverse real-world scenarios. The investigation includes an analysis of prompt engineering strategies, such as hand-crafted prompts, prompt ensembles, and prompt learning techniques. We introduce a vision-text-space ensemble that significantly boosts the average performance compared to a text-space-only ensemble. Additionally, our comparative study delves into leveraging existing test-time adaptation methods originally designed for image classification tasks. Experimental evaluations conducted across various datasets and model architectures demonstrate the efficacy of different adaptation strategies. We further give insights into the importance of updating the vision encoder and whether it is beneficial to update the text encoder. Code is available at: https://github.com/mariodoebler/test-time-adaptation
@@ 1 Introduction
2 Related Work
2.1 Foundation Models
2.2 Online Test-time Adaptation
Online test-time adaptation adapts the model to an unknown domain shift directly during inference, leveraging all available test samples. One successful line of work recalculates the batch normalization (BN) statistics during test-time [41] to mitigate covariate shift caused by corruption. Although updating only the BN statistics is computationally efficient, it has its limitations, especially regarding natural domain shifts. As a result, recent TTA methods additionally incorporate model weight updates through self-training. TENT [46], for example, showcased that minimizing the entropy with respect to the batch normalization parameters can successfully improve the performance for single-target adaptation. Building upon this idea, EATA [33] introduces a loss weighting and filtering scheme that accounts for the reliability and diversity of a sample. Furthermore, they use elastic weight consolidation [21] to mitigate catastrophic forgetting [30] on the initial training domain. However, accessing data from the initial training domain may not always be feasible in practical scenarios. To prevent a model from collapsing to a trivial solutions induced by confidence maximization, [26, 31] apply diversity regularizers. Other works, such as [5, 9] also employ contrastive learning to mitigate a domain shift during test-time.
Recent research has tackled even more challenging scenarios, such as dealing with temporally correlated data. LAME [3] focuses on adapting the model’s output using Laplacian adjusted maximum-likelihood estimation. On the other hand, NOTE [14], RoTTA [51], and DAB [11] introduce a buffer to simulate an i.i.d. stream. To handle large and noisy gradients that can promote model collapse, SAR [34] proposes a sharpness-aware and reliable entropy minimization method. Building upon SAR, DeYO [23] incorporates a confidence metric that measures the extent to which the probability of pseudo-label decreases after applying an image transformation that distorts the shape of the objects.
In the work of [29], recent TTA methods are evaluated on a broad range of possible TTA scenarios, termed Universal TTA. Their proposed method ROID [29] puts emphasis on using certainty and diversity weighting to prevent model collapse during adaptation and introduces weight ensembling at test-time to prevent the adapted model from overfitting on the current distribution shift. CMF [24] builds upon the framework of ROID and replaces weight ensembling by continual momentum filtering. It utilizes the Kalman filtering algorithm to find a source model that is robust against catastrophic forgetting while mainting high flexibility.
Due to their multimodality and zero-shot generalization capabilities, VLMs offer new possibilities for TTA. One approach that focuses on adapting the prompt space is test-time prompt tuning (TPT) [42]. It is loosely based on the supervised context optimization (CoOp) [52]. The basic idea of TPT is to optimize the prompt context via entropy minimization of an augmented batch based on a single sample. In this work we want to give new perspectives on how to deal with VLMs, namely CLIP, in the context of online test-time adaptation.
3 Prompts and Vision-space Ensembles
@@
Vision-language foundation models such as CLIP [37] aim to learn a joint embedding space for the vision and language modality. This is achieved by aligning the representations of images and their associated textual descriptions through contrastive learning. To extract the embeddings, CLIP leverages a separate encoder for the vision and text modality, denoted here as $f_{\mathrm{vision}}$ and $f_{\mathrm{text}}$, respectively. After successfully training the encoders on typically millions of image-text pairs, the learned joint embedding space allows to associate similar concepts across modalities, resulting in cross-modal understanding.
In the case of zero-shot classification with a hand-crafted prompt, the procedure involves the following steps. Let $\bm{x}_{t}\in\mathbb{R}^{H\times W\times C}$ be the current test image at time step $t$ with height $H$, width $W$, and $C$ channels, and $\bm{z}_{t}=f_{\mathrm{vision}}(\bm{x}_{t})$ denote its corresponding representation. In addition, let $\{\bm{t}_{1},\bm{t}_{2},\dots,\bm{t}_{K}\}$ be a textual representation for each of the $K$ classes, obtained by embedding short phrases (templates) like ”a photo of a {classname}.” into the text embedding space. Now, to determine the class label for an image, its representation $\bm{z}_{t}$ is first paired with each of the $K$ text representations $(\bm{t}_{k},\bm{z}_{t})$. Then, the cosine similarity $s_{k}=\mathrm{sim}(\bm{t}_{k},\bm{z}_{t})$ is computed for each pair. The final model prediction simply corresponds to the class with the highest similarity score or highest softmax probability. The latter can be computed with
where $\tau$ is a temperature. It’s worth noting that the text embeddings $\{\bm{t}_{1},\bm{t}_{2},\dots,\bm{t}_{K}\}$ are typically precomputed once before inference, ensuring efficiency during prediction.
3.1 Prompt Engineering
While using a simple phrase like a photo of a {classname}. can already work exceptionally well, the performance of VL models heavily depends on the utilized prompt and its encoded representation [52]. Thus, writing better hand-crafted prompts can significantly improve the performance of the model. For instance, in the context of ImageNet-R [17] and a ViT-B-16 model pretrained by OpenAI, employing the prompt template ”depiction of a {classname}.” reduces the error rate from 26.0% to 23.6%. Similarly, for datasets like EuroSAT [15] that contain low-resolution satellite photos, using a prompt such as a blurry satellite photo of {classname}. decreases the error rate from 58.5% to 46.3%. These examples underscore the importance of well-designed prompts to optimize model performance.
Instead of relying on a single prompt template, Radford et al. [37] also proposed to use a list of $J$ different templates. An example can look like the following list [”a photo of a {classname}.”, ”a sketch of a {classname}.”, ”a painting of a {classname}.”]. By averaging the text representations obtained from all templates for class $k$, i.e.,
$\bar{\bm{t}}_{k}=\frac{1}{J}\sum_{j=1}^{J}\bm{t}_{kj},$ (2)
While the ensemble approach described earlier uses a predefined list of hand-crafted prompt templates, CuPL [36] introduces a novel strategy that harnesses the power of a large language model to generate a class-specific prompt list. Specifically, the LLM is asked to write descriptive sentences that encapsulate the discriminative features of the various classes. In the case of the category goldfish, the prompt list might look like [”Most goldfish have a shiny gold or orange color.”, ”A goldfish in a bowl.”, …]. These descriptive prompts help to improve the performance of the VL model without requiring any expert knowledge.
3.2 Learning Prompts
Zhou et al. [52] introduced context optimization to offline fine-tune CLIP-like vision-language models with a few labeled training examples $\{\bm{x}_{i},\bm{y}_{i}\}_{i=1}^{N}$, where $\bm{y}_{i}\in\mathbb{R}^{K}$ is the one-hot encoded category of image $\bm{x}_{i}$. Unlike before, where the context of the prompt (such as ”a photo of a”) was either fixed or manually tuned, it is now learnable. This involves representing the context with a few learnable token embeddings and then minimizing, for example, a cross-entropy (CE) loss according to
where []delimited-[][$\cdot$][ ⋅ ] is the Iverson bracket and $\beta$ is a threshold. After the context is updated one (or several) times, regular zero-shot classification can be performed. While [42] demonstrate the efficacy of this approach, the computational overhead is huge, since for each test image, there are 64 forward passes through the image encoder and at least 2 forward passes through the text encoder - one to learn a better context and another to acquire the new text representations.
3.3 Vision-Text-Space Ensemble
The effectiveness of methods like TPT depends on the creation of suitable training examples via test-time augmentation, which can be identified with confidence-based filtering, for example. Feng et al. [13] take this approach one step further by additionally exploiting Stable Diffusion [39] to generate new images that resemble the current test image. Since leveraging a diverse set of augmented test samples might also be helpful for the previous hand-crafted or LLM-based prompt ensemble approaches, we introduce a Vision-Text-Space Ensemble (VTE), as depicted in Fig. 1, which creates an ensemble in both spaces. Following TPT, VTE utilizes the same test-time augmentation strategy and entropy-based confidence filtering to extract reliable samples from the artificially generated batch. The representations of the identified samples are then averaged via the equation
$\bar{\bm{z}}_{t}=\frac{1}{\rho B}\sum_{i=1}^{B}[e_{ti}\leq\beta]\bm{z}_{ti},$ (5)
where $\bar{\bm{z}}_{t}$ is subsequently utilized to compute the similarity scores according to $s_{k}=\mathrm{sim}(\bar{\bm{t}}_{k},\bar{\bm{z}}_{t})$. Note that this procedure is again optimization-free and, unlike TPT, does not require any forward passes through the text encoder during test-time.
3.4 Experiments
@@
3.4.1 Datasets, Models, and Metric
We follow the continual test-time adaptation setting in [29] and evaluate the models’ robustness on ImageNet (validation set) and its variants. ImageNet-C [16] includes 15 types of corruptions with 5 severity levels applied to the validation images of ImageNet (IN). For the natural domain shifts, we consider ImageNet-R [17], ImageNet-Sketch [47], as well as ImageNet-D109, a variation of ImageNet-D [40] introduced in [29]. While ImageNet-R contains 30,000 examples depicting different renditions of 200 IN classes, ImageNet-Sketch contains 50 sketches for each of the 1,000 IN classes. Additionally, we report results for ImageNet-V2 [38] and ImageNet-A [18]. ImageNet-V2 is an independent test set containing 10,000 natural images covering all 1,000 IN classes. ImageNet-A comprises 7,500 adversarial examples for a subset of 200 IN categories.
To evaluate categories outside the ImageNet context, we follow [42] and report results for ten datasets, covering fine-grained classifications including species of plants or animals (Flowers102 [32], OxfordPets [35]), scenes (SUN397 [50]), textures (DTD [7]), food (Food101 [2]), transportation (StanfordCars [22], Aircraft [27]), human actions (UCF101 [43]), satellite images (EuroSAT [15]), and general objects (Caltech101 [12]).
We utilize six different architectures for our evaluation, encompassing CLIP with a ResNet-50 (RN50) backbone, as well as ViT [10] models of various sizes: ViT-B-16, ViT-L-14, and ViT-H-14, all pretrained by OpenAI. Furthermore, we consider EVA-02-B-16 and EVA02-L-14 from [44]. Our evaluation metric is based on the error rate.
3.4.2 Results
The results of the diverse prompt-based methods are illustrated in Table 1. Here, Source denotes employing zero-shot classification with different prompt strategies: utilizing the single prompt ”a photo of a {classname}.”, an ensemble of hand-crafted prompt templates following [37], the CuPL [36] prompts generated by an LLM, and a combination of both ensemble and CuPL prompts referred to as ”All Prompts”.
As shown in Table 1, all methods substantially improve on the single prompt baseline. While the LLM generated prompts of CuPL outperform the hand-crafted ensemble on five out of seven ImageNet variations for both architectures, even better results can be achieved by leveraging all prompts. This even outperforms the optimization based approach of TPT on four out of seven IN variations, while requiring only a fraction of its computational effort, i.e., one image forward versus 64 image forwards, 2 text forwards, and one backward. However, the best results are achieved by our VTE approach, which significantly outperforms all other baselines. Although this comes at the cost of an increased computational complexity compared to the hand-crafted approaches, VTE is still faster than TPT during inference, since it does not require any text forwards or backwards. We also find, that the performance of VTE can be further improved by employing a better prompt list. For ViT-B-16, for example, using All Prompts decreases the average error from $38.6\%$ to $38.0\%$.

### v2

@@ Abstract
In deep learning, maintaining model robustness against distribution shifts is critical. This work explores a broad range of possibilities to adapt vision-language foundation models at test-time, with a particular emphasis on CLIP [37] and its variants. The study systematically examines prompt-based techniques and existing test-time adaptation methods, aiming to improve the robustness under distribution shift in diverse real-world scenarios. Specifically, the investigation covers various prompt engineering strategies, including handcrafted prompts, prompt ensembles, and prompt learning techniques. Additionally, we introduce a vision-text-space ensemble that substantially enhances average performance compared to text-space-only ensembles. Since online test-time adaptation has shown to be effective to mitigate performance drops under distribution shift, the study extends its scope to evaluate the effectiveness of existing test-time adaptation methods that were originally designed for vision-only classification models. Through extensive experimental evaluations conducted across multiple datasets and diverse model architectures, the research demonstrates the effectiveness of these adaptation strategies. Code is available at: https://github.com/mariodoebler/test-time-adaptation
Keywords:
test-time adaptation vision-language models
@@ 1 Introduction
2 Related Work
2.1 Foundation Models
2.2 Online Test-Time Adaptation
Online test-time adaptation adapts the model to an unknown domain shift directly during inference, leveraging all available test samples. One successful line of work recalculates the batch normalization (BN) statistics during test-time [41] to mitigate covariate shift caused by corruption. Although updating only the BN statistics is computationally efficient, it has its limitations, especially regarding natural domain shifts. As a result, recent TTA methods additionally incorporate model weight updates through self-training. TENT [46], for example, showcased that minimizing the entropy with respect to the batch normalization parameters can successfully improve the performance for single-target adaptation. Building upon this idea, EATA [33] introduces a loss weighting and filtering scheme that accounts for the reliability and diversity of a sample. Furthermore, they use elastic weight consolidation [21] to mitigate catastrophic forgetting [30] on the initial training domain. However, accessing data from the initial training domain may not always be feasible in practical scenarios. To prevent a model from collapsing to trivial solutions induced by confidence maximization, [26, 31] apply diversity regularizers. Other works, such as [5, 9] also employ contrastive learning to mitigate a domain shift during test-time.
Recent research has tackled even more challenging scenarios, such as dealing with temporally correlated data. LAME [3] focuses on adapting the model’s output using Laplacian adjusted maximum-likelihood estimation. On the other hand, NOTE [14], RoTTA [51], and DAB [11] introduce a buffer to simulate an i.i.d. test stream. To handle large and noisy gradients that can promote trivial solutions, SAR [34] proposes a sharpness-aware and reliable entropy minimization method. Building upon SAR, DeYO [24] incorporates a confidence metric that measures the extent to which the probability of pseudo-label decreases after applying an image transformation that distorts the shape of the objects.
In the work of [29], recent TTA methods are evaluated on a broad range of possible TTA scenarios, termed Universal TTA. Their proposed method ROID [29] puts emphasis on using certainty and diversity weighting to prevent the occurrence of trivial solutions during the adaptation. To further preserve the model’s generalization capabilities and overcome catastrophic forgetting, ROID introduces weight ensembling. This approach continuously combines the weights of initial source model with those of the current adaptation model during test-time. CMF [23] builds upon the ROID framework and replaces weight ensembling by continual momentum filtering. It utilizes a Kalman filter to derive a model that is both resilient to catastrophic forgetting and highly adaptable.
Due to their multimodality and zero-shot generalization capabilities, VLMs offer new possibilities for TTA. One approach that focuses on adapting the prompt space is TPT [42]. It is inspired by the supervised context optimization (CoOp) [52] approach, but differs by directly optimizing the prompt context during test-time. This is achieved by minimizing the entropy of an augmented batch generated from a single test sample. In this work, we aim to provide new perspectives on how to deal with VLMs, namely CLIP, in the context of online test-time adaptation.
3 Prompts and Vision-Text-Space Ensembles
In this chapter, we first revisit the underlying principles of vision-language foundation models such as CLIP, along with their approach to perform zero-shot classification. We then introduce improved prompting strategies, including our novel approach, and provide a comprehensive benchmark of all methods. Building upon this foundation, Chapter 4 explores the combination of improved prompting techniques with existing TTA methods that update the model parameters.
Vision-language foundation models, in particular CLIP [37], aim to learn a joint embedding space for the vision and language modality. This is achieved by aligning the representations of images and their associated textual descriptions through contrastive learning. To extract the embeddings, CLIP leverages a separate encoder for the vision and text modality, denoted here as $f_{\mathrm{vision}}$ and $f_{\mathrm{text}}$, respectively. After successfully training the encoders on typically hundreds of millions of image-text pairs, the learned joint embedding space allows to associate similar concepts across modalities, resulting in cross-modal understanding.
To perform zero-shot classification with a hand-crafted prompt, the procedure involves the following steps. Let $\bm{x}_{t}\in\mathbb{R}^{H\times W\times C}$ be the current test image at time step $t$ with height $H$, width $W$, and $C$ channels, and $\bm{z}_{t}=f_{\mathrm{vision}}(\bm{x}_{t})$ denote its corresponding representation. In addition, let $\{\bm{t}_{k}\}_{k=1}^{K}$ be a textual representation for each of the $K$ classes, obtained by embedding short phrases (templates) like "a photo of a {classname}." into the text embedding space. Now, to determine the class label for an image, its representation $\bm{z}_{t}$ is first paired with each of the $K$ text representations $(\bm{t}_{k},\bm{z}_{t})$. Then, the cosine similarity $s_{k}=\mathrm{sim}(\bm{t}_{k},\bm{z}_{t})$ is computed for each pair. The final model prediction simply corresponds to the class with the highest similarity score or highest softmax probability. The latter can be computed with
where $\tau$ is a temperature. Note that the text embeddings $\{\bm{t}_{k}\}_{k=1}^{K}$ are typically precomputed once before inference, ensuring efficiency during test-time.
3.1 Prompt Engineering
While using a simple phrase like a photo of a {classname}. can already work exceptionally well, the performance of VL models heavily depends on the utilized prompt and its encoded representation [52]. Thus, writing better hand-crafted prompts can significantly improve the performance of the model. For instance, in the context of ImageNet-R [16] and a ViT-B-16 model pretrained by OpenAI, employing the prompt template "depiction of a {classname}." reduces the error rate from 26.0% to 23.6%. Similarly, for datasets like EuroSAT [15] that contain low-resolution satellite photos, using a prompt such as "a blurry satellite photo of {classname}." decreases the error rate from 58.5% to 46.3%. These examples underscore the importance of well-designed prompts to maximize performance.
Instead of relying on a single prompt template, Radford et al. [37] also proposed to use a list of $J$ different templates. An example can look like the following list ["a photo of a {classname}.", "a sketch of a {classname}.", "a painting of a {classname}."]. By averaging the text representations obtained from all templates for class $k$, i.e.,
$\bar{\bm{t}}_{k}=\frac{1}{J}\sum_{j=1}^{J}\bm{t}_{kj},$ (2)
While the ensemble approach described earlier uses a predefined list of hand-crafted prompt templates, CuPL [36] introduces a novel strategy that harnesses the power of a large language model to generate a class-specific prompt list. Specifically, the LLM is asked to write descriptive sentences that encapsulate the discriminative features of the various classes. In the case of the category goldfish, the prompt list might look like ["Most goldfish have a shiny gold or orange color.", "A goldfish in a bowl.", …]. These descriptive prompts help to improve the performance of the VL model without requiring any expert knowledge.
3.2 Learning Prompts
Zhou et al. [52] introduced context optimization to offline fine-tune CLIP-like vision-language models with a few labeled training examples $\{\bm{x}_{i},\bm{y}_{i}\}_{i=1}^{N}$, where $\bm{y}_{i}\in\mathbb{R}^{K}$ is the one-hot encoded category of image $\bm{x}_{i}$. Unlike before, where the context of the prompt (such as "a photo of a") was either fixed or manually tuned, it is now learnable. This involves representing the context with a few learnable token embeddings, which are then optimized by minimizing, for example, a cross-entropy (CE) loss according to
where []delimited-[][$\cdot$][ ⋅ ] is the Iverson bracket and $\beta$ is a threshold. After the context is updated one (or several) times, regular zero-shot classification can be performed. While [42] demonstrate the effectiveness of this approach, it incurs substantial computational overhead. Specifically, each test image results in 64 forward passes through the image encoder and at least 2 forward passes through the text encoder - one for learning an improved context and another to acquire the new text representations.
3.3 Vision-Text-Space Ensemble
@@
The effectiveness of methods like TPT depends on the creation of suitable training examples via test-time augmentation, which can subsequently be identified with confidence-based filtering, for example. Feng et al. [13] take this approach one step further by additionally exploiting Stable Diffusion [39] to generate new images that resemble the current test image. Since leveraging a diverse set of augmented test samples might also be helpful for the previous hand-crafted or LLM-based prompt ensemble approaches, we introduce a Vision-Text-Space Ensemble (VTE), which creates an ensemble in both spaces. Following TPT, VTE utilizes the same test-time augmentation strategy and entropy-based confidence filtering to extract reliable samples from the artificially generated batch. The representations of the identified samples are then averaged via the equation
$\bar{\bm{z}}_{t}=\frac{1}{\rho B}\sum_{i=1}^{B}[e_{ti}\leq\beta]\bm{z}_{ti},$ (5)
where $\bar{\bm{z}}_{t}$ is subsequently utilized to compute the similarity scores according to $s_{k}=\mathrm{sim}(\bar{\bm{t}}_{k},\bar{\bm{z}}_{t})$. Note that this procedure is again optimization-free and, unlike TPT, does not require any forward passes through the text encoder during test-time. An illustration of VTE is also shown in Fig. 1.
3.4 Experiments
3.4.1 Datasets, Models, and Metric
We follow the continual test-time adaptation setting in [29] and evaluate the models’ robustness on ImageNet (validation set) and its variants. ImageNet-C [17] includes 15 types of corruptions with 5 severity levels applied to the validation images of ImageNet (IN). For the natural domain shifts, we consider ImageNet-R [16], ImageNet-Sketch [47], as well as ImageNet-D109, a variation of ImageNet-D [40] introduced in [29]. While ImageNet-R contains 30,000 examples depicting different renditions of 200 IN classes, ImageNet-Sketch contains 50 sketches for each of the 1,000 IN classes. Additionally, we report results for ImageNet-V2 [38] and ImageNet-A [18]. ImageNet-V2 is an independent test set containing 10,000 natural images covering all 1,000 IN classes. ImageNet-A comprises 7,500 adversarial examples for a subset of 200 IN classes.
To evaluate categories outside the ImageNet context, we follow [42] and report results for ten datasets, covering fine-grained classifications including species of plants or animals (Flowers102 [32], OxfordPets [35]), scenes
(SUN397 [50]), textures (DTD [7]), food (Food101 [2]), transportation (StanfordCars [22], Aircraft [27]), human actions (UCF101 [43]), satellite images (EuroSAT [15]), and general objects (Caltech101 [12]).
While the main experiments are conducted using CLIP with a ViT-B-16 and ViT-L-14 backbone [10], we later also explore additional architectures, including a ResNet-50 (RN50) and a ViT-H-14 model, all pretrained by OpenAI. Furthermore, we consider EVA-02-B-16 and EVA02-L-14 from [44]. Our evaluation metric is based on the error rate.
3.4.2 Results
@@
The results of the diverse prompt-based methods are illustrated in Table 1. Here, Source denotes employing zero-shot classification with different prompt strategies: utilizing the single prompt "a photo of a {classname}.", an ensemble of hand-crafted prompt templates following [37], the CuPL [36] prompts generated by an LLM, and a combination of both ensemble and CuPL prompts referred to as "All Prompts".
As shown in Table 1, all methods substantially improve on the single prompt baseline. While the LLM generated prompts of CuPL outperform the hand-crafted ensemble on five out of seven ImageNet variations for both architectures, better results can be achieved by leveraging all prompts. This even outperforms the optimization based approach TPT on four out of seven IN variations, while requiring only a fraction of its computational effort, i.e., one image forward versus 64 image forwards, 2 text forwards, and one backward. However, the best results are achieved by our VTE approach, which significantly outperforms all other baselines. Although this comes at the cost of an increased computational complexity compared to the hand-crafted approaches, VTE is still faster than TPT during inference due to not requiring any backwards or text forwards through the respective encoder. We also find, that the performance of VTE can be further improved by employing a better prompt list. For ViT-B-16, for example, using All Prompts decreases the average error from $38.6\%$ to $38.0\%$.
@@ 4 Updating Model Parameters with Test-Time Adaptation
While the focus of the previous section has been on leveraging prompts and vision-text-space ensembles, in this section we want to put emphasis on a surprisingly underexplored topic, namely leveraging existing TTA methods for adapting vision-language models. The idea is straightforward, the text encoder of a CLIP model is frozen, allowing to precompute the text embeddings $\{\bm{t}_{k}\}_{k=1}^{K}$, ensuring efficiency during prediction. Given an image embedding $\bm{z}_{t}$, the cosine similarity can be computed $s_{t}=\mathrm{sim}(\bm{t}_{k},\bm{z}_{t})$. Treating the cosine similarities as the network’s logits, the output probabilites can be received through Eq. 1. In this way we can treat any CLIP model as a common image classifier, enabling the application of any existing TTA method for image classification. Note that it is also possible to update the text encoder’s parameters, but for now, we limit our analysis to only updating the parameters of the image encoder. In the following experiments, a batch size of 64 test samples per time step $t$ is employed.
4.1 Test-Time Normalization for CLIP
First, we investigate the performance of BN–1, a common procedure in TTA, which recalculates the batch normalization (BN) statistics using the current test batch. While Schneider et al. [41] showed that recalculating the batch normalization statistics during test-time can significantly reduce the error rate for models pretrained on ImageNet, we investigate whether this is also the case for a CLIP model that was trained on millions of data samples covering a much broader data distribution. In Figure 3 the zero-shot performance (source) and BN–1 performance is illustrated for CLIP with a RN50 and RN101 backbone using a single prompt. It can be clearly seen that, unlike for models pretrained on ImageNet, the average performance across the investigated datasets substantially decreases when applying BN–1. For RN50 the average error rate increases from 50.2% to 74.1% and for RN101 from 46.8% to 71.0%. This can be possibly attributed to much larger batch sizes and a much broader data distribution used during CLIP pretraining. A similar phenomenon is described in [29], where employing BN–1 for a regular ImageNet pretrained RN50 decreases the error rate on ImageNet-C from 82.0% to 68.6% in a continual TTA setting, but increases to 82.5% in a mixed-domains TTA setting, where all corruptions of ImageNet-C are randomly suffled within the test sequence. Since BN–1 is employed by most TTA methods during adaptation, we conclude that RN backbones are not feasible. Instead we focus our following analysis on vision transformers that do not employ BN.